A farmer sows a certain crop. It costs $250,000 to buy the seed, prepare the ground, and sow the crop. In one year's time it will cost $95,000 to harvest the crop. If the crop will be worth $350,000, and the interest rate is 10%, what is the net present value (NPV) of this investment? Select one: A. $240,000 B. $567,103 C. $18,181 D. $87,103
Answer #1

NPV = present value of cash inflows-present value of cash outflows

Present value (PV)of Cash inflows:

Interest rate =10%

350,000 we will get in one year time

So PV of cash inflows = 350,000*1/1.1=318181

PV of Cash outflows:

250,000 costs at the beginning hence no discounting is required

95,000 will get in one year time hence discounting required

PV of cash outflows = 95000/1.1 = 86363

NPV = 318181-250000-86363

NPV =-18181 (Option C)